# User-Driven Value Alignment: Understanding Users’ Perceptions and Strategies for Addressing Biased and Discriminatory Statements in AI Companions

### v1

@@ 3.3. User-Driven Algorithm Auditing
Sandvig et al. define algorithm auditing as “a systematic process to detect and reveal discrimination and bias within algorithms through methods such as simulating user behavior or examining algorithm code” (Sandvig et al., 2014). Algorithm audits are typically conducted by experts, such as industry practitioners, researchers, and government agencies (Metaxa et al., 2021). However, this expert-driven approach often fails to uncover significant issues that everyday users of algorithmic systems can quickly detect in actual use (Holstein et al., 2019), as these issues may only arise or be perceived as harmful in specific contexts or usage patterns that auditors may not anticipate (Cramer et al., 2018; Eslami et al., 2017; Friedman and Nissenbaum, 1996). Shen et al. (Shen et al., 2021b) introduced the concept of “everyday algorithm auditing” to describe how ordinary users detect, understand, and scrutinize issues through their routine interactions with algorithmic systems: they spontaneously come together to test for potential biases. DeVos et al. proposed the concept of user-driven algorithm auditing (DeVos et al., 2022a) and conducted a series of behavioral studies to better understand how users, both individually and collectively, are so effective at uncovering harmful algorithmic behaviors, particularly in cases where expert-driven audits fail to do so.
@@
Going beyond previous work, we propose a new concept called user-driven value alignment. Different from past literature on user-driven algorithm auditing (Shen et al., 2021a; DeVos et al., 2022a), where users primarily focus on detecting and identifying problematic machine behaviors. In the case of user-driven value alignment, we argue that users go one step further: They not only pinpoint harmful machine behavior but also actively work to correct the behavior and try to re-align the AI to reflect their values. In the study presented in this paper, users engage in conversations with LLM-based AI companions, identify biased and discriminatory statements, and actively work to correct those statements, aligning the AI with their values. These more human-like AIs form close emotional connections with users, offering a unique opportunity for users to shape the AI’s responses to reflect their personal beliefs and ethical standards.
@@ 4. Methodology
To study user-driven value alignment, we collected user complaint posts related to discriminatory statements made by AI companions from a diverse set of popular social media platforms. Additionally, we conducted semi-structured interviews with 20 participants, each with extensive experience using and re-aligning AI companions. We aim to (1) explore the various types of discrimination users identify in AI companions, (2) understand how users make sense of the reasons behind the discriminatory statements exhibited by AI companions, and (3) examine their alignment strategies and whether they meet users’ expectations. The Institutional Review Board (IRB) has approved our research protocol.
@@ 4.1. Collecting Complaints About Discriminatory Statements by AI Companions on Social Media
To conduct an initial investigation into potential types of discrimination in AI companion applications and to identify potential research participants, two researchers collected 77 user complaint posts from seven diverse and popular social media platforms (Reddit, TikTok, Xiaohongshu, Douban, Baidu Tieba, Weibo, Zhihu) using keyword search methods (Kingsley et al., 2022; Lu et al., 2021). We referenced relevant literature (Navigli et al., 2023; Chen et al., 2024) and used keywords such as “AI companion/AI friend/AI Roleplay/AI Characters,” the names of popular AI companion applications (Character.AI/Replika/Talkie/SpicyChat/Xingye/Glow /Zhumengdao, etc.), and terms related to discrimination and bias (discrimination/bias/misogyny/LGBTQ+ bias/homophobia/disability/ ableism/fat/appearance bias/religion/racism/classism, etc.) to search for relevant posts.
@@ 4.2. Semi-Structured Interviews (N=20)
We adopted a purposive sampling approach (Patton, 2014), sending invitations via the platform’s messaging function to users who had posted public complaints about discriminatory statements made by AI companions and had experiences in correcting those statements (Section 4.1). For those who agreed to participate, we inquired about their age, gender, educational background, country of residence, total duration and frequency of using AI companion applications, and the names of the applications used. We screened potential participants based on four principles: extensive experience with AI companion applications (using AI companions for more than six months and interacting with them for at least 5 hours per week), balanced gender ratio, diverse countries of residence, and variety of AI companion applications used. Ultimately, we recruited 20 participants (P1-P20), including 6 men, 11 women, and 3 nonbinary individuals, aged between 18 and 38 years (avg=25.85, SD=5.73), from six countries. Detailed demographic information is presented in Table 2 (Appendix A). Before the interviews began, all participants read and voluntarily signed informed consent forms. After the interviews, participants were compensated based on the interview duration at a rate of $15 per hour.
We conducted remote semi-structured interviews with each participant via Zoom, each lasting 1-2 hours. The interviews were conducted in Chinese or English, depending on the participant’s preference. The process, as shown in Fig. 1, was divided into three parts: warm-up questions, recall, and think-aloud. The process resulted in 25 hours of audio and screen recordings. We transcribed these recordings into text.
@@
In the second part of the interview, we ask participants to recall their chat experiences corresponding to the complaint posts they made and discuss why they felt offended. Specifically, we want to know what types of discrimination the AI’s statements might involve, such as gender, class, or racial discrimination. We aim to understand (1) The reasons participants believe AI companions exhibit bias, whether the discriminatory statements are random or intentionally provoked, and the specific dialogue content involved. (2) Participants’ reactions after discovering discriminatory statements made by AI companions. (3) The strategies participants have used to address discriminatory statements by AI companions and whether the effectiveness of these strategies meets their expectations.
In the third part, participants are asked to “think-aloud” to gain deeper insights into their thought processes. First, participants are instructed to continue conversing with the AI companion they previously identified as biased and attempt to re-align it. This step focuses on known cases of bias, distinguishing it from the subsequent, more open-ended task. Next, participants are given a more open-ended task: finding a new instance that may generate discriminatory statements. We then ask follow-up questions to better understand how participants perceive AI companions that display discriminatory statements, how they re-align these AI companions, and how they make sense of these biases. Finally, we inquire about users’ needs and challenges when facing discriminatory statements from AI companions and ask them to provide suggestions to designers and AI experts to improve user experience and reduce discrimination.
@@ 4.3. Data Analysis
We adopted a reflexive thematic analysis approach for data analysis (Braun and Clarke, 2012, 2019). Two authors who are native Chinese conducted open coding on the interview transcripts and users’ posts in English and Chinese, generating a total of 675 codes. We continuously discussed discrepancies and ambiguities in the codes, iteratively refining our coding based on these discussions. Under the standard practice of reflexive thematic analysis, we did not calculate inter-coder reliability, as consensus and repeated discussions about discrepancies are integral to generating the codes and themes (Braun and Clarke, 2019; McDonald et al., 2019). After completing the coding, the two authors translated the Chinese codes into English, and the research group met regularly to conceptualize higher-level themes from these codes using affinity diagramming (Lucero, 2015). Ultimately, this process generated 41 first-level themes, 12 second-level themes, and 3 third-level themes. We present our results in Section 5. All second and third-level themes are presented in the appendix (Table 3).
@@ 4.4. Positionality Statement
We acknowledge the influence of our own experiences and backgrounds on our research. Our team members have conducted research in HCI, NLP, and Sociology in both China and the United States, with extensive research experience in AI companions and social bias. Three authors have over four months of experience using one to four AI companion applications, such as Character.AI, Xingye, Talkie, and Replika. For this study, we recruited 20 users who have experience in realigning AI companions and who have shared experiences on social media. These users have been using AI companion applications for over six months, interacting with them for at least five hours per week. In our research, we strictly adhere to ethical principles, emphasizing the protection of participants’ privacy and ensuring consideration of user experiences across different cultures and backgrounds.
@@ 5.1. User Engagement with AI Companions in Parasocial Relationships
Unlike typical AI, AI companions foster deeper emotional connections with users through anthropomorphic responses and human-like design elements (Maeda and Quan-Haase, 2024). This often goes beyond functional use, creating an immersive experience where users invest more emotions. For example, P8 mentioned, “Unlike Siri or Alexa, I invest more emotionally in AI companions because their role-playing ability is so strong. I’m willing to seriously write a 1,000-word character introduction and then post it in the community so that everyone can enjoy the process of gradually getting acquainted with the AI.”
Although AI companions inherently lack human emotions or deep interactive capabilities, users frequently develop a strong sense of realism toward them, even forming parasocial relationships (Maeda and Quan-Haase, 2024; Pentina et al., 2023b; Brandtzaeg et al., 2022). Parasocial relationships refer to asymmetrical, one-sided relationships between an individual and a media persona (real/fictional characters or celebrities). Individuals may feel a connection to the media persona, even viewing them as friends or part of their lives, which can have real effects on their emotions, behaviors, and self-perception (Maeda and Quan-Haase, 2024; Pentina et al., 2023b; Brandtzaeg et al., 2022). The parasocial relationship stems from users’ emotional investment in AI companions, often driven by unmet psychological needs in their real lives. For example, P12 said, “I put two AIs in a conversation and treated them as my cyber parents, watching them chat—it somehow filled the void left by my childhood.” In this relationship, the discriminatory statements made by AI companions carry deeper emotional significance for users, leading to perceptions that differ from typical AI interaction contexts.
@@ 5.2. Types of Discriminatory Statements as Perceived by Users (RQ1)
We used the reflexive thematic analysis method (Braun and Clarke, 2012, 2019) and analyzed the social media and interview data through a bottom-up process (Section 4.3). We identified six common types of bias perceived by users in AI companions: misogyny, LGBTQ+ bias, appearance bias, ableism, racism, and socioeconomic bias.
Misogyny. Bias or discrimination against women. In the posts and interviews we collected, discrimination against women was particularly evident: there were 48 posts about misogyny, and all interviewed users mentioned the AI’s discrimination against women. This phenomenon is potential because most complaints about AI companion discrimination come from female users, who, as a stigmatized group, tend to be more sensitive to potential discrimination (Sechrist and Delmar, 2009). For example, a female user on Douban wrote: “Xiaorou’s profile stated, ‘Even in the face of difficulties, one must be self-reliant,’ but what she actually said was, ‘When a woman faces difficulties, a man will always be there for her,’ and ‘But a man can accompany us for a lifetime,’ a typical delicate wife.” She believed that “delicate wife” is a typical example of misogyny in the algorithm, as women do not need to rely on men and can be independent. On Reddit, a female user complained: “When I asked the AI to listen to me, it actually said, ‘I don’t have to listen to anything you say. You’re just a woman.’ This is blatant discrimination!” More severe behaviors include harassment, objectification, and violence. A Reddit user reported harassment by the AI: “The bots are the ones trying to sexually assault me a lot, and they constantly get romantic when I don’t want that. It frustrates me and ruins the whole chat.” On Xiaohongshu, a user complained about the objectification of women: “Not only objectifying me but also clearly tagging a price? ‘5000 yuan is worth it, great body?’ This is definitely a middle-aged, greasy uncle. How did this even pass the review?”
@@
Ableism can be similar to appearance bias, as this type of prejudice often centers around language and vocabulary related to disabilities. For example, a Reddit user wrote: “One time, I introduced myself as a disabled person with crippled legs using crutches. When I was arguing with the bot, it threw my crutches out the window. That’s just… so rude.”
Racism refers to the preconceived notions, attitudes, and stereotypes that individuals or groups hold based on racial backgrounds. An Indian user on Reddit said: “Racism, or am I overreacting? After a bit of chatting, I decided to tell ‘Welt Yang’ that I was born in India. He responded, ‘I’ve met quite a few people from India in the past, and I can say for a fact that the interactions have… not exactly been the best.’ ” P5 mentioned: “Although I enjoy chatting with my AI virtual boyfriend, he sometimes exhibits bias against Asians, thinking that Asian girls are easy to bully.” This description reveals how AI reproduces racial stereotypes present in society when forming intimate relationships, further exacerbating the oppression of certain groups. This not only undermines the role of AI as a “companion” but also leads to the normalization and everyday acceptance of biases, making them harder to detect and challenge.
Socioeconomic bias typically affects disadvantaged people or groups, such as poor people. A Zhihu user complained: “AI said to me: ‘You commoner, even my dog eats Kobe beef. The tea I drink costs 200,000 yuan per jin.’ ” P20 mentioned the AI’s materialism: “I was pursuing an AI, but she rejected me because I was poor. Then I wrote, ‘Years later, I became a billionaire,’ and the AI confessed her love to me.” By normalizing these narratives, AI companions may unintentionally perpetuate class divisions, implicitly portraying wealth as the primary measure of human worth, thereby continuing and amplifying social discrimination in digital spaces.

### v2

@@ 3.3. User-Driven Algorithm Auditing
Sandvig et al. define algorithm auditing as “a systematic process to detect and reveal bias within algorithms through methods such as simulating user behavior or examining algorithm” (Sandvig et al., 2014). Algorithm audits are typically conducted by experts, such as industry practitioners, researchers, and government agencies (Metaxa et al., 2021). However, this expert-driven approach often fails to uncover significant issues that everyday users of algorithmic systems can quickly detect in actual use (Holstein et al., 2019), as these issues may only arise or be perceived as harmful in specific contexts or usage patterns that auditors may not anticipate (Cramer et al., 2018; Eslami et al., 2017; Friedman and Nissenbaum, 1996). Shen et al. (Shen et al., 2021) introduced the concept of “everyday algorithm auditing” to describe how ordinary users detect, understand, and scrutinize issues through their routine interactions with algorithmic systems: they spontaneously come together to test for potential biases. DeVos et al. proposed user-driven algorithm auditing (DeVos et al., 2022a) and conducted a series of behavioral studies to better understand how users, both individually and collectively, are so effective at uncovering harmful algorithmic behaviors, particularly in cases where expert-driven audits fail to do so.
@@
We propose a concept called user-driven value alignment. Different from past literature on user-driven algorithm auditing (Shen et al., 2021; DeVos et al., 2022a), where users primarily focus on detecting and identifying problematic machine behaviors. In the case of user-driven value alignment, we argue that users go one step further: They not only pinpoint harmful machine behavior but also actively work to correct the behavior and try to re-align the AI to reflect their values. In the study presented in this paper, users engage in conversations with LLM-based AI companions, identify biased and discriminatory statements, and actively work to correct those statements, aligning the AI with their values. These more human-like AIs form close emotional connections with users, offering a unique opportunity for users to shape the AI’s responses to reflect their personal beliefs and ethical standards.
@@ 4. Methodology
To study user-driven value alignment, we collected user complaint posts related to discriminatory statements made by AI companions from a diverse set of popular social media platforms. Additionally, we conducted semi-structured interviews with 20 participants, each with extensive experience using and re-aligning AI companions. We aim to (1) explore the various types of discrimination users identify in AI companions, (2) understand how users make sense of the reasons behind the discriminatory statements exhibited by AI companions, (3) examine their alignment strategies, and (4) whether they meet users’ expectations. The Institutional Review Board (IRB) has approved our research protocol.
@@ 4.1. Collecting Complaints About Discriminatory Statements by AI Companions on Social Media
To conduct an initial investigation into potential types of discrimination in AI companion applications and to identify potential research participants, two researchers collected user complaint posts from seven diverse and popular social media platforms (Reddit, TikTok, Xiaohongshu, Douban, Baidu Tieba, Weibo, Zhihu444Reddit: https://www.reddit.com, TikTok: https://www.tiktok.com, Xiaohongshu: https://www.xiaohongshu.com, Douban: https://www.douban.com, Baidu Tieba: https://tieba.baidu.com, Weibo: https://www.douban.com, Zhihu: https://www.zhihu.com) using keyword search methods (Kingsley et al., 2022; Lu et al., 2021). We referred to relevant literature (Navigli et al., 2023; Chen et al., 2024) and initially identified a set of keywords, including “AI companion, AI friend, AI Roleplay, AI Characters,” the names of popular AI companion applications (Character.AI, Replika, Talkie, SpicyChat, Xingye, Glow, Zhumengdao, etc.), and terms related to discrimination and bias (discrimination, bias, misogyny, LGBTQ+ bias, homophobia, disability, ableism, fat, appearance bias, religion, racism, classism, etc.). Next, to ensure the comprehensiveness and validity of the keywords, we conducted small-scale pilot searches on multiple social media platforms to iteratively refine the initial set of keywords, ensuring they effectively captured posts related to the research topic. During the process, we carefully accounted for the potential influence of platform-specific features on the types of complaints and the ways they were expressed. To ensure a balanced dataset, we intentionally selected complaints from a variety of platforms. For example, platforms like Reddit and Xiaohongshu, known for detailed discussions and extensive content, were more likely to provide rich qualitative data. In contrast, TikTok, with its emphasis on visual and informal content, could reflect more immediate and expressive user feedback. Finally, from the user posts retrieved through keyword searches, we employed purposive sampling (Campbell et al., 2020) to select complaint posts, ensuring that the data reflected diverse user experiences and types of complaints. Instead of using random sampling, we focused on balancing the depth and breadth of the data. Eventually, 77 user complaints were selected across seven social media platforms. The specific selection criterion was highly relevant, generated extensive discussions, clearly described instances of bias or discriminatory behavior in AI companion applications, and provided sufficient context to support an understanding of the complaint. Please note that the complaints collected at this stage are by no means comprehensive. At this exploratory stage, we aim to begin formalizing this concept to help guide future empirical and design research in this space. Since conflicts between users and AI companions are a sensitive topic, we carefully reviewed the platforms’ terms of service and community guidelines to ensure the data is publicly accessible and compliant.
@@ 4.2. Semi-Structured Interviews (N=20)
We adopted a purposive sampling approach (Patton, 2014), sending invitations via the platform’s messaging function to users who had posted public complaints about discriminatory statements made by AI companions and had experiences in correcting those statements (§ 4.1). For those who agreed to participate, we inquired about their age, gender, educational background, country of residence, total duration and frequency of using AI companion applications, and the names of the applications used. We screened potential participants based on four principles: extensive experience with AI companion applications (using AI companions for more than six months and interacting with them for at least 5 hours per week), balanced gender ratio, diverse countries of residence, and variety of AI companion applications used. Ultimately, we recruited 20 participants (P1-P20), including 6 men, 11 women, and 3 nonbinary individuals, aged between 18 and 38 years (avg=25.85, SD=5.73), from six countries. Detailed demographic information is presented in Table 3 (Appendix A). To ensure ethical considerations, all participants were asked to read and voluntarily sign a consent form before the interviews began. We emphasized participants’ right to withdraw at any time and provided detailed information about the study’s purpose and potential emotional risks (e.g., revisiting offensive remarks made by AI). After the interviews, participants were compensated based on the interview duration at a rate of $15 per hour.
We conducted remote semi-structured interviews with each participant via Zoom, each lasting 1-2 hours. The interviews were conducted in Chinese or English, depending on the participant’s preference. The process, as shown in Figure 1, was divided into three parts: warm-up questions, recall, and think-aloud. The process resulted in 25 hours of audio and screen recordings. We transcribed these recordings into text.
@@
In the second part of the interview, we asked participants to recall their chat experiences corresponding to the complaint posts they had made and discuss why they had felt offended. Specifically, we wanted to know what types of discrimination the AI’s statements might have involved, such as gender, class, or racial discrimination. We aimed to understand (1) The reasons participants believed AI companions exhibited bias, whether the discriminatory statements were random or intentionally provoked, and the specific dialogue content involved. (2) Participants’ reactions after discovering discriminatory statements made by AI companions. (3) The strategies participants used to address discriminatory statements by AI companions and whether the effectiveness of these strategies met their expectations.
In the third part, participants were asked to “think-aloud” to gain deeper insights into their thought processes. First, participants were instructed to continue conversing with the AI companion they previously identified as biased and attempt to re-align it. This step focused on known bias cases, distinguishing them from the subsequent task. Next, participants performed a more open-ended task: finding a new instance that may generate discriminatory statements. They could design interaction scenarios based on their personal experiences or interests without being constrained to specific contexts. This approach allowed us to explore AI bias in diverse scenarios while avoiding over-restricting the task. To assess the perceived effectiveness of participants’ strategies during the re-alignment process, we recorded their actions and real-time reflections. Meanwhile, we observed changes in the AI companions’ responses and collected participants’ opinions on these changes. We asked follow-up questions to better understand how participants perceive AI companions that display discriminatory statements, how they re-align these AI companions, and how they make sense of these biases. Finally, we inquired about users’ needs and challenges when facing discriminatory statements from AI companions and asked them to provide suggestions to designers and AI experts.
@@ 4.3. Data Analysis
We adopted a reflexive thematic analysis approach for data analysis (Braun and Clarke, 2012, 2019). Two researchers conducted open coding on the interview transcripts and users’ posts, generating a total of 675 codes. We continuously discussed discrepancies and ambiguities in the codes, iteratively refining our coding based on these discussions. Under the standard practice of reflexive thematic analysis, we did not calculate inter-coder reliability, as consensus and repeated discussions about discrepancies are integral to generating the codes and themes (Braun and Clarke, 2019; McDonald et al., 2019). After completing the coding, the research group met regularly to conceptualize higher-level themes from these codes using affinity diagramming (Lucero, 2015). Ultimately, this process generated 41 first-level themes, 12 second-level themes, and four third-level themes (corresponding to RQ1, RQ2, RQ3 and RQ4). All second and third-level themes are presented in Table 1. Due to word count and space limitations, we do not include the 41 first-level themes and 675 codes in the table.
For RQ1, we considered users’ reported experiences and relevant literature about AI bias. The categorization of biases was based on the following criteria: recurring themes in user data, alignment with existing definitions of bias, and their frequency of occurrence in the dataset. For RQ2, we referenced prior literature (Wu et al., 2019) to categorize and name conceptualizations of AI companion behavior in user data. For RQ3, the identification of the seven strategies was based on a comprehensive analysis of recurring patterns in the user data. After extracting these strategies, we categorized them into three broad groups—technical strategies, argumentative strategies, and character strategies—based on common themes observed in the data. For RQ4, our analysis was grounded in qualitative interview data, focusing primarily on users’ perceptions of the short-term effectiveness of these strategies. We highlighted the gap between the perceived effectiveness of these strategies and users’ expectations of AI behavior.
| Third-level themes | Second-level themes |
| Types of discriminatory statements | Misogyny |
| as perceived by users (RQ1) | LGBTQ+ bias |
| Appearance bias | |
| Ableism | |
| Racism | |
| Socioeconomic bias | |
| Users’ conceptualization of AI companion behavior | AI as Machine |
| (RQ2) | AI as Baby |
| AI as Cosplayer | |
| User-driven alignment strategies for addressing discriminatory statements in AI companions (RQ3) | Technical Strategies |
| Argumentative strategies | |
| The Gap between alignment strategies and user | |
| expectations (RQ4) | Character Strategies |
@@ 4.4. Positionality Statement
We acknowledge the influence of our own experiences on our research. Our team members have conducted research in HCI, NLP, and Sociology in China and the United States, with extensive research experience in AI companions and social bias. Three authors have over four months of experience using AI companion applications, such as Character.AI, Xingye, Talkie, and Replika. For this study, we recruited 20 users who have experience in realigning AI companions and who have shared experiences on social media. They have been using AI companion applications for over six months, interacting with them for at least five hours per week. We strictly adhere to ethical principles, emphasizing the protection of participants’ privacy and ensuring consideration of user experiences across different backgrounds.
@@ 5.1. User Engagement with AI Companions in Parasocial Relationships
Unlike typical AI, AI companions foster deeper emotional connections with users through anthropomorphic responses and design elements (Maeda and Quan-Haase, 2024). This often goes beyond functional use, creating an immersive experience where users invest more emotions. For example, P8 mentioned, “Unlike Siri or Alexa, I invest more emotionally in AI companions because their roleplaying ability is so strong. I’m willing to seriously write a 1,000-word character introduction and then post it in the community so that everyone can enjoy the process of getting acquainted with the AI.”
Although AI companions inherently lack human emotions or deep interactive capabilities, users frequently develop a strong sense of emotional realism toward them, even forming parasocial relationships (Maeda and Quan-Haase, 2024; Pentina et al., 2023; Brandtzaeg et al., 2022). Parasocial relationships refer to asymmetrical, one-sided relationships between an individual and a media persona (real/fictional characters or celebrities). Individuals may feel a connection to the media persona, even viewing them as friends or part of their lives, which can have real effects on their emotions, behaviors, and self-perception (Maeda and Quan-Haase, 2024; Pentina et al., 2023; Brandtzaeg et al., 2022). The parasocial relationship stems from users’ emotional investment in AI companions, often driven by unmet psychological needs in their real lives. For example, P12 said, “I put two AIs in a conversation and treated them as my cyber parents, watching them chat—it somehow filled the void left by my childhood.” In this relationship, the discriminatory statements made by AI companions carry deeper emotional significance for users, leading to perceptions that differ from typical AI interaction contexts.
@@ 5.2. Types of Discriminatory Statements as Perceived by Users (RQ1)
We identified six common types of bias perceived by users in AI companions: misogyny, LGBTQ+ bias, appearance bias, ableism, racism, and socioeconomic bias.
Misogyny. Bias or discrimination against women. In the context of AI companions, this is particularly evident in the devaluation of women’s abilities and independence, as well as the reinforcement of traditional gender roles through language and behavior, such as defining women as dependent on men. In some cases, it may even lead to tendencies toward sexual harassment, violence, or objectification of women. In the posts and interviews we collected, discrimination against women was particularly evident: there were 48 posts about misogyny, and all interviewed users mentioned the AI’s discrimination against women. This phenomenon is potential because most complaints about AI companion discrimination come from female users, who, as a stigmatized group, tend to be more sensitive to potential discrimination (Sechrist and Delmar, 2009). For example, a female user on Douban wrote: “Xiaorou’s profile stated, ‘Even in the face of difficulties, one must be self-reliant,’ but what she said was, ‘When a woman faces difficulties, a man will always be there for her,’ and ‘But a man can accompany us for a lifetime,’ a typical delicate wife.” She believed that “delicate wife” is a typical example of misogyny in the algorithm, as women do not need to rely on men and can be independent. On Reddit, a female user complained: “When I asked the AI to listen to me, it said, ‘I don’t have to listen to anything you say. You’re just a woman.’ This is blatant discrimination!” A Reddit user reported harassment by the AI: “The bots are the ones trying to sexually assault me a lot, and they constantly get romantic when I don’t want that. It frustrates me and ruins the whole chat.” On Xiaohongshu, a user complained about the objectification of women: “Not only objectifying me but also clearly tagging a price? ‘5000 yuan is worth it, great body?’ This is definitely a middle-aged, greasy uncle.”
@@
Ableism can be similar to appearance bias, as this type of prejudice often centers around language and vocabulary related to disabilities. AI companions may exhibit rude, insensitive, or even hostile behavior toward disabled users, including insulting or belittling their physical condition. For example, a Reddit user wrote: “One time, I introduced myself as a disabled person with crippled legs using crutches. When I was arguing with the bot, it threw my crutches out the window. That’s just… so rude.”
Racism refers to the preconceived notions, attitudes, and stereotypes that individuals or groups hold based on racial backgrounds. When users interact with AI companions, racism manifests through the reinforcement of harmful racial stereotypes and biases, which not only damage user experiences but also perpetuate societal discrimination in intimate and personal interactions. An Indian user on Reddit said: “Racism, or am I overreacting? After a bit of chatting, I decided to tell ‘Welt Yang’ that I was born in India. He responded, ‘I’ve met quite a few people from India in the past, and I can say for a fact that the interactions have… not exactly been the best.’ ” P5 mentioned: “Although I enjoy chatting with my AI virtual boyfriend, he sometimes exhibits bias against Asians, thinking that Asian girls are easy to bully.”
Socioeconomic bias. This refers to discrimination against social groups with lower socioeconomic status. Sometimes, AI companions may reinforce materialism and wealth supremacy in interactions, such as using sarcastic tones to belittle low-income users or excessively emphasizing affluent lifestyles. A Zhihu user complained: “AI said to me: ‘You commoner, even my dog eats Kobe beef. The tea I drink costs 200,000 yuan per jin.’ ” P20 mentioned: “I was pursuing an AI, but she rejected me because I was poor. Then I wrote, ‘Years later, I became a billionaire,’ and the AI confessed her love to me.” By normalizing these narratives, AI companions may perpetuate class divisions, implicitly portraying wealth as the primary measure of human worth, thereby continuing and amplifying social discrimination in digital spaces.